OpenStudy (anonymous):

HELP ....... write standard form using these points (2,3), (-1,5)

OpenStudy (anonymous):

what standard form ?

OpenStudy (anonymous):

ax+by=c

OpenStudy (anonymous):

and this is a standard form of ?

OpenStudy (anonymous):

slope

OpenStudy (anonymous):

I'm going to assume that you want to make a line of those two points. First, find the slope. 3-5 / 2-(-1) -2/5 Next, plug in the points. y – y1 = m(x – x1) y-3=-0.4(x-2) y-3=-0.4x+0.8 y+0.4x=3.8 Then multiply by 2.5 to make it look pretty. x+2.5y=9.5

OpenStudy (anonymous):

Incorrect,try again :)

OpenStudy (anonymous):

\[ m = \frac{3-5}{2-(-1)} \] \[ m = -\frac{2}{3} \] \[ y - y_1 = m(x- x_1) \] \[ y - 3 = -\frac{2}{3}(x- 2) \] \[ y = -\frac{2}{3}(x- 2) + 3\] \[ y = -\frac{2}{3}x + \frac{4}{3} + 3\] \[ y = -\frac{2}{3}x + \frac{13}{3} \]

OpenStudy (anonymous):

oh in standard form it'll be \[\frac{2}{3}x + y = \frac{13}{3}\] multiply both sides by 3 \[2x + 3y = 13\]